# Towards Neural Scaling Laws for Time Series Foundation Models

### v1

@@ Abstract
Scaling laws offer valuable insights into the design of time series foundation models (TSFMs). However, previous research has largely focused on the scaling laws of TSFMs for in-distribution (ID) data, leaving their out-of-distribution (OOD) scaling behavior and the influence of model architectures less explored. In this work, we examine two common TSFM architectures—encoder-only and decoder-only Transformers—and investigate their scaling behavior on both ID and OOD data. These models are trained and evaluated across varying parameter counts, compute budgets, and dataset sizes. Our experiments reveal that the log-likelihood loss of TSFMs exhibits similar scaling behavior in both OOD and ID settings. We further compare the scaling properties across different architectures, incorporating two state-of-the-art TSFMs as case studies, showing that model architecture plays a significant role in scaling. The encoder-only Transformers demonstrate better scalability than the decoder-only Transformers, while the architectural enhancements in the two advanced TSFMs primarily improve ID performance but reduce OOD scalability. While scaling up TSFMs is expected to drive performance breakthroughs, the lack of a comprehensive understanding of TSFM scaling laws has hindered the development of a robust framework to guide model scaling. We fill this gap in this work by synthesizing our findings and providing practical guidelines for designing and scaling larger TSFMs with enhanced model capabilities.
@@ 1 Introduction
The neural scaling law quantitatively describes how model performance grows with the scaling of three basic training factors: model parameters, computational resources and training dataset size. Establishing such scaling laws is crucial for developing TSFMs, as it provides a framework for predicting expected performance gains, enabling the community to rationally allocate efforts toward key designs. The exploration on scaling laws for TSFMs is still in an initial stage; recent research has primarily focused on studying ID scaling behavior (Edwards et al., 2024; Shi et al., 2024a). In practical applications, TSFMs primarily face challenges from unseen scenarios(Wang et al., 2024), where OOD forecasting capability is most critical. This raises an unresolved question: do neural scaling laws also apply to predict out-of-distribution forecasting performance? Moreover, various architectures of TSFMs have been arising, but they typically focus on performance improvement at specific scales. No studies have investigated the scaling behaviors across different architectures, leaving a key question unanswered: how do model architectures affect scalability? Although we are seeing an increasing investment in training resources for TSFMs, the bottlenecks and potential driving factors for developing larger TSFMs remain unclear. This raises another practical question: how to design TSFMs from the perspective of scalability?
In this paper, we aim to provide empirical answers to the above research questions. To investigate the scaling laws in OOD scenarios, we trained a family of encoder-only Transformer-based TSFMs, varying three basic training factors: model sizes, compute budgets, and training set sizes. We evaluated their performance on both ID and OOD test sets and established scaling laws for three training factors in each scenario. To examine the impact of model architecture on scaling behavior, we trained decoder-only Transformer based TSFMs and compared them with the encoder-only versions. Additionally, we included two state-of-the-art TSFMs, Moirai (Woo et al., 2024) and Chronos (Ansari et al., 2024), as case studies for detailed analysis. Our experiment results suggest that the log-likelihood loss of TSFMs exhibits similar scaling behavior in both OOD and ID scenarios; encoder-only Transformer has better scalability than decoder-only Transformer; the architectural modifications introduced by two advanced TSFMs mainly improve ID performance but compromise OOD scalability. Based on the findings and comparative analysis, we finally provided design principles for TSFMs from a scaling perspective.
@@
  • Scaling laws across model architectures. We investigate the scaling patterns of different architectures of TSFMs, offering some insights for scalable TSFM design.
@@ 2 Preliminary
To investigate the scaling laws of TSFMs, we curated a large, diverse, and balanced dataset for pre-training. Leveraging this dataset, we trained both encoder-only and decoder-only transformers and two state-of-the-art TSFMs: Chronos and Moirai. For comparative analysis, we evaluated these models on (i) in-distribution and (ii) out-of-distribution test sets, focusing on key performance metrics to examine the scaling behavior across architectures.
@@ 2.1 Datasets
A large scale, diverse, balanced and high quality pre-training dataset is the foundation to build FMs. To this end, we constructed our time series corpus for TSFM pre-training from the large-scale open time series archive, Lotsa (Woo et al., 2024). The corpus comprises approximately $15$B time points from $39$ datasets spanning seven distinct domains. To ensure that the model performs fairly across all domains, we maintained a balanced ratio of data from different domains. Furthermore, we performed quality filtering on the corpus by constraining the signal-to-noise ratio of a time series to be greater than $20$ dB, ensuring that the pre-training corpus exhibits strong predictability. A detailed breakdown of the data sources is provided in Appendix A, with a summary in Table 1.
@@
Domain Transport Climate Energy CloudOps Health Sales Web Sum
@@
Time points 4.82B 4.73B 2.34B 2.15B 240M 140M 600M 14.46B
Proportion 33.31% 32.71% 16.15% 14.86% 1.61% 0.96% 0.40% 100%
To assess the impact of pre-training data scale on model performance, we partitioned the corpus into subsets containing $10$M, $100$M, and $1$B time points, ensuring that each subset maintained similar diversity. For each subset, $95$% of the data was allocated for model training, with the remaining $5$% reserved as a validation set to evaluate in-distribution forecasting performance. Additionally, we used a subset from a widely recognized long-sequence prediction benchmark (Wu et al., 2023) to test the model’s out-of-distribution forecasting capabilities. This subset includes test data from the ETTh1-2, ETTm1-2, electricity, and weather datasets.
Figure 1: Architectures of Baseline Time Series Foundation Models. The the most widely used two Transformer architectures, encoder-only Transformer and decoder-only Transformer, are selected to be our baselines. The time series is divided into patches, each treated as a token input to the Transformer model. Shaded patches represent the future horizon to be predicted.
@@ 2.2 Models
TSFMs are predominantly built upon the Transformer architecture (Wen et al., 2022). For our baseline models, we selected two widely adopted architectures: the encoder-only Transformers and the decoder-only Transformers. We also include the state-of-the-art models from both architectures, Moirai (Woo et al., 2024) (encoder-only) and Chronos (Ansari et al., 2024) (decoder-only), in our study for comparison. Although Chronos-T5 is technically an encoder-decoder Transformer, we categorize it as analogous to decoder-only Transformers because it primarily relies on the decoder’s auto-regressive method for prediction. The primary distinction between encoder-only and decoder-only architectures lies in the attention mechanisms applied to the inputs, as illustrated in Figure 1. To better adapt them for time series forecasting, we introduce three key modifications in input layer, positional encoding and prediction head. More details are given in appendix B.
@@
Rotary Position Embedding. This technique (RoPE) has rapidly gained popularity as a positional encoding method in recent large language models (Su et al., 2024). Given its proven effectiveness in improving time series forecasting performance (Woo et al., 2023), we adopt RoPE as a replacement for the original Transformer’s positional encoding. RoPE encodes absolute positions using a rotation matrix while embedding relative position dependencies directly into the self-attention mechanism.
Mixture of Distributions. Our models are designed to predict the probability distribution of future time series. However, real-world time series often exhibit complex distributions, including outliers, heavy tails, and extreme skew, which pose significant challenges for accurate modeling. To address these complexities, we incorporate a more flexible output likelihood by utilizing Student-t mixture models (Flunkert et al., 2017). Compared to the commonly used Gaussian mixture models, Student-t mixture models offer greater robustness in handling outliers and heavy-tailed distributions.
Our models are characterized by several key hyperparameters: the number of layers ($n_{\rm layer}$), the input/output dimensions of the residual stream ($d_{\rm m}$), the dimensions of the intermediate feed-forward layers ($d_{\rm ff}$), the number of attention heads per layer ($n_{\rm heads}$), and the dimension of the attention output ($d_{\rm head}$). The overall model size can be expressed as:
@@
where the embedding layer, prediction head, biases and other sub-leading terms are excluded for a cleaner scaling laws. The embedding layer uses a patch size of 32 with $32{d_{\rm m}}$ parameters. The mixture distribution prediction head comprises multiple independent linear layers that predict the Student-t mixture distribution parameters for each patch, with $512{d_{\rm m}}$ parameters in total. In the study, we explore models with $\sim 10^{3}$ to $\sim 10^{8}$ trainable parameters.
@@ 2.3 Training and Evaluation Details
In this study, we focus exclusively on univariate time series forecasting to avoid the confounding effects introduced by multivariate time series, such as variable interactions, correlations, and the complexities of modeling multivariate relationships. Future research will address these factors, aiming to establish more comprehensive scaling laws for multivariate time series models.
Training Details. Our training objective is to optimize the mixture distribution log-likelihood. We utilize the AdamW optimizer with a batch size of 128 and a cosine learning rate scheduler with a linear warm-up of $10^{4}$ training steps, training for a total of $10^{5}$ steps. To facilitate learning data representations across diverse domains with varying series lengths and sample sizes, we visited each sample with probability $p_{i}=t_{i}/T$, where $t_{i}$ is the series’ time points and $T$ is the corpus’ total time points. We then randomly selected a segment from each chosen sample.
Evaluation Details. We evaluate the model on a randomly selected $10$% subset of the test data every $10^{3}$ steps to reduce computational costs. For performance measurement, we observed that non-normalized metrics like MAE and MSE are highly sensitive to the amplitude of time series data, often causing the overall average to be disproportionately influenced by high-amplitude datasets. To mitigate this issue, we utilize the normalized metric, mean absolute percentage error (MAPE), along with the log-likelihood loss, to assess forecasting performance. Detailed descriptions of these metrics are provided in the Appendix C.3.
@@ 3 Scaling Laws for Time Series Foundation Models
In this section, we first present experimental results using the encoder-only Transformer to explore scaling laws across different data distributions. Following this, we conduct a comparative study on the scaling behavior of encoder-only and decoder-only TSFMs, Moirai and Chronos, to investigate how model architectures influence the scalability of time series models.
@@ 3.1 Scaling Laws Across Data Distributions
Figure 2: Parameter Scaling. The scaling effect of trainable model parameters on the in-distribution (ID) and out-of-distribution (OOD) forecasting performance, which is evaluated using log-likelihood and MAPE metrics. When evaluated with log-likelihood, both ID and OOD results follow an approximate power law scaling with parameter count, exhibiting consistent trends across different data distributions.
Parameter Scaling. In Figure 2, we display the ID and OOD performance of a wide variety of encoder-only Transformers, ranging from small models with 3K parameters to large models with 300M parameters. We trained models on the full 15B pre-training corpus to convergence and report the minimum log-likelihood loss and MAPE. We can see that both ID and OOD performance roughly follow power-law behavior over five orders of magnitude in model sizes. Formally, the power law can be expressed as:
@@
where $L$ is the performance metric function (i.e., MAPE, or log-likelihood), $N$ is a given parameter count, $N_{c}$ is the normalization coefficient, and $\alpha_{N}$ is the exponent value that indicates the degree of performance improvement expected as we scale up $N$.
Observing the log-likelihood metric, the lines fitting the scaling laws for both ID and OOD data exhibits a roughly constant shift and close slopes. This implies that while models incur a consistent performance bias when transferred to OOD data, their scaling patterns correlate well with their performance on the ID data. When evaluated using MAPE, the power-law for OOD scenario shows a bigger exponent value than ID scenario. This indicates that increasing model size yields greater improvements in OOD performance than ID performance. In other words, for models with weak OOD generalization capabilities, increasing model size may enable them to perform equally well on both ID and OOD data.
Figure 3: Compute Scaling. The computation scaling results show that model performance approximately follows a power law trend with increasing compute, consistent across both ID and OOD scenarios. The ID and OOD results demonstrate that a lower bound exists for the loss and MAPE under a given computational budget.
Compute Scaling. Following the similar method in (Kaplan et al., 2020), we estimate the compute budget using the formula $C=6NBS$, where $B$ is the batch size, $S$ is the number of parameter updates, i.e. the input sequence length, and 6 is the factor to account for the forward and backward passes. The ID and OOD test loss for compute budget varying over six orders of magnitude are shown in Fig 3. We see that the optimal results for each compute budget are achieved by different model sizes $N$, but the lowest loss decreases according to a approximate power law with respect to the amount of training compute. The lowest losses appear as the heavy lines, which can be fit with
@@
We observed significant noise in log-likelihood and MAPE during training, which may be caused by the learning rate scheduler and the random sampling evaluation strategy. Similar to parameter scaling, the ID and OOD log-likelihood shows similar scaling patterns; however, when evaluated using MAPE, the OOD power law shows larger exponent values than the ID power law.
Figure 4: Data Scaling. The blue and red plots illustrate how data volume affects the ID and OOD forecasting performance of encoder-only Transformers, evaluated using log-likelihood and MAPE metrics. The results indicate that in both scenarios, model performance scales approximately as a power law with increasing data volume.
Data Scaling. We display empirical trends for the performance as function of dataset size $D$ in Fig 4. For the trend, we trained multiple 1B encoder-only Transformers on a series of subsets of the 15B pre-training dataset and report the averaged evaluation results during training. We see that the log-likelihood and MAPE can be fit well with simple power-law
@@
Different from the parameter scaling and computational scaling, we found that when evaluated using log-likelihood, ID and OOD performance do not exhibit the same scaling behavior. Instead, ID performance is more sensitive to the scaling of dataset size compared to OOD performance. However, similar to the observations on the other two factors, the impact of data scaling on MAPE in OOD data is greater than its impact on MAPE in ID data. This suggests that the scaling of various factors yields greater improvements in OOD performance than ID performance.
Cross-Distribution Scaling Effects Summary. We summarize the key findings on how model ID and OOD performance scales with the model parameters, data volume, and compute. (1) In both ID and OOD univariate time series forecasting, model performance follows a simple power law as a function of model parameters, data volume, and compute. (2) For log-likelihood loss, the model exhibits similar scaling patterns in both ID and OOD scenarios, in terms of model size or compute resources. (3) When using MAPE as the metric, scaling of all three factors results in greater improvements in OOD performance compared to ID performance.
Figure 5: Scaling Laws of Encoder-only vs. Decoder-only Transformer. This figure presents a comparison of scaling behaviors on log-likelihood between encoder-only and decoder-only Transformer across three different axes: number of parameters, compute, and dataset size. Overall, encoder-only Transformer shows better scalability on model parameters, computation and dataset sizes across ID and OOD data.

### v2

@@ Abstract
Scaling laws offer valuable insights into the design of time series foundation models (TSFMs). However, previous research has largely focused on the scaling laws of TSFMs for in-distribution (ID) data, leaving their out-of-distribution (OOD) scaling behavior and the influence of model architectures less explored. In this work, we examine two common TSFM architectures—encoder-only and decoder-only Transformers—and investigate their scaling behavior on both ID and OOD data. These models are trained and evaluated across varying parameter counts, compute budgets, and dataset sizes. Our experiments reveal that the negative log-likelihood of TSFMs exhibits similar scaling behavior in both OOD and ID settings. We further compare the scaling properties across different architectures, incorporating two state-of-the-art TSFMs as case studies, showing that model architecture plays a significant role in scaling. The encoder-only Transformers demonstrate better scalability than the decoder-only Transformers in ID data, while the architectural enhancements in the two advanced TSFMs primarily improve ID performance but reduce OOD scalability. While scaling up TSFMs is expected to drive performance breakthroughs, the lack of a comprehensive understanding of TSFM scaling laws has hindered the development of a robust framework to guide model scaling. We fill this gap in this work by synthesizing our findings and providing practical guidelines for designing and scaling larger TSFMs with enhanced model capabilities.
@@ 1 Introduction
The neural scaling law (Kaplan et al., 2020) quantitatively describes how model performance grows with the scaling of three basic training factors: model parameters, computational resources and training dataset size. Establishing such scaling laws is crucial for developing TSFMs, as it provides a framework for predicting expected performance gains, enabling the community to rationally allocate efforts toward key designs. The exploration on scaling laws for TSFMs is still in an initial stage; recent research has primarily focused on studying ID scaling behavior (Edwards et al., 2024; Shi et al., 2024a). In practical applications, TSFMs primarily face challenges from unseen scenarios (Wang et al., 2024), where OOD forecasting capability is most critical. This raises an unresolved question: do neural scaling laws also apply to predict out-of-distribution forecasting performance? Moreover, various architectures of TSFMs have been arising, but they typically focus on performance improvement at specific scales. No studies have investigated the scaling behaviors across different architectures, leaving a key question unanswered: how do model architectures affect scalability? Although we are seeing an increasing investment in training resources for TSFMs, the bottlenecks and potential driving factors for developing larger TSFMs remain unclear. This raises another practical question: how to design TSFMs from the perspective of scalability?
In this paper, we aim to provide empirical answers to the above research questions. To investigate the scaling laws in OOD scenarios, we trained a family of encoder-only Transformer-based TSFMs, varying three basic training factors: model sizes, compute budgets, and training set sizes. We evaluated their performance on both ID and OOD test sets and established scaling laws for three training factors in each scenario. To examine the impact of model architecture on scaling behavior, we trained decoder-only Transformer based TSFMs and compared them with the encoder-only versions. Additionally, we included two state-of-the-art TSFMs, Moirai (Woo et al., 2024) and Chronos (Ansari et al., 2024), as case studies for detailed analysis. Our experiment results suggest that the negative log-likelihood loss of TSFMs exhibits similar scaling behavior in both OOD and ID scenarios; encoder-only Transformers have a similar scalability with decoder-only Transformers, with a slight advantage in ID data; the architectural modifications introduced by two advanced TSFMs mainly improve ID performance but compromise OOD scalability. Based on the findings and comparative analysis, we finally provided design principles for TSFMs from a scaling perspective.
@@
  • Scaling laws across model architectures. We investigate the scaling patterns of different TSFM architectures, showing the scalability varies depending on the model architecture and the design of specific modules.
@@ 2 Preliminary
To investigate the scaling laws of TSFMs, we curated a large, diverse, and balanced dataset for pre-training. Leveraging this dataset, we trained both encoder-only and decoder-only transformers within two state-of-the-art TSFMs: Moirai and Chronos. For comparative analysis, we evaluated these models on (i) in-distribution and (ii) out-of-distribution test sets, focusing on key performance metrics to examine the scaling behavior across architectures.
@@ 2.1 Datasets
A large scale, diverse, balanced and high quality pre-training dataset is the foundation to build FMs. To this end, we constructed our time series corpus for TSFM pre-training from the large-scale open time series archive, LOTSA (Woo et al., 2024). The corpus comprises approximately $17$B time points from $39$ datasets spanning seven distinct domains. To ensure that the model performs fairly across all domains, we maintained a balanced ratio of data from different domains. Furthermore, we performed quality filtering on the corpus by constraining the signal-to-noise ratio of a time series to be greater than $20$ dB, ensuring that the pre-training corpus exhibits strong predictability. A detailed breakdown of the data sources is provided in Appendix A, with a summary in Table 1.
@@
Domain Transport Climate Energy CloudOps Health Sales Web Total
@@
Time points 4.82B 4.73B 4.76B 2.15B 232M 140M 40M 16.8B
Proportion 28.52% 28.06% 28.21% 12.76% 1.38% 0.83% 0.24 % 100%
To assess the impact of pre-training data scale on model performance, we partitioned the corpus into subsets containing $10$M, $100$M, and $1$B time points, ensuring that each subset maintained similar diversity. For each subset, $95$% of the data was allocated for model training, with the remaining $5$% reserved as a validation set to evaluate in-distribution forecasting performance. Additionally, we used a subset from a widely recognized long-sequence prediction benchmark (Wu et al., 2023) to test the model’s out-of-distribution forecasting capabilities. To further enhance the reliability, we also incorporated a subset of the Monash dataset (Godahewa et al., 2021) as additional OOD test data. The details of the dataset composition and properties are provided in Appendix A Table LABEL:tab:ood_dataset.
Figure 1: Architectures of Baseline Time Series Foundation Models. As the most widely used two Transformer architectures, encoder-only Transformer and decoder-only Transformer are selected to as our baseline. A time series is divided into multiple patches, each treated as a token and fed into the Transformer model. The shaded patches represent the future horizon to be predicted.
@@ 2.2 Models
TSFMs are predominantly built upon the Transformer architecture (Wen et al., 2022). For our baseline models, we selected two widely adopted architectures: the encoder-only Transformers (Woo et al., 2024) (Moirai) and the decoder-only Transformers (Ansari et al., 2024) (Chronos). The primary distinction between them lies in the attention mechanisms applied to the inputs, as illustrated in Figure 1. To better adapt them for time series forecasting, we introduce three key modifications in input layer, positional encoding and prediction head. More details are given in Appendix B.
@@
Rotary Position Embedding. This technique (RoPE) has rapidly gained popularity as a positional encoding method in recent large language models (Su et al., 2024). Given the improved pre-training efficiency observed with RoPE (Woo et al., 2023), we adopt RoPE as a replacement for the original Transformer’s positional encoding. RoPE encodes absolute positions using a rotation matrix while embedding relative position dependencies directly into the self-attention mechanism.
Mixture of Distributions. Our models are designed to predict the probability distribution of future time series. However, real-world time series often exhibit complex distributions, including outliers, heavy tails, and extreme skew, which pose significant challenges for accurate modeling. To address these complexities, we incorporate a more flexible output likelihood by utilizing Student-T mixture models (Flunkert et al., 2017). Compared to the commonly used Gaussian mixture models, Student-T mixture models offer greater robustness in handling outliers and heavy-tailed distributions. An empirical comparison between the two mixture distributions is shown in Appendidx B Figure 9.
Our models are characterized by several key hyper-parameters: the number of layers ($n_{\rm layer}$), the input/output dimensions of the residual stream ($d_{\rm m}$), the dimensions of the intermediate feed-forward layers ($d_{\rm ff}$), the number of attention heads per layer ($n_{\rm heads}$), and the dimension of the attention output ($d_{\rm head}$). The overall model size can be expressed as:
@@
where the embedding layer, prediction head, biases and other sub-leading terms are excluded for a cleaner scaling laws. The embedding layer uses a patch size of 32 with $32{d_{\rm m}}$ parameters. The mixture distribution prediction head comprises multiple independent linear layers that predict each Student-t mixture distribution parameter for a patch separately, with $512{d_{\rm m}}$ parameters in total. In the study, we explore models with $\sim 10^{3}$ to $\sim 10^{8}$ trainable parameters.
@@ 2.3 Training and Evaluation Details
In this study, we focus exclusively on uni-variate time series forecasting to avoid the confounding effects introduced by multivariate time series, such as variable interactions, correlations, and the complexities of modeling multivariate relationships. Future research will address these factors, aiming to establish more comprehensive scaling laws for multivariate time series models.
Training Details. Our training objective is to optimize the mixture distribution log-likelihood. We utilize the AdamW optimizer with a batch size of 128, and a maximum learning rate of $10^{-3}$ with a linear warm-up of $10^{4}$ training steps, followed by cosine decay for the remaining $9\times 10^{4}$ steps. To facilitate learning data representations across diverse domains with varying series lengths and sample sizes, we visited each sample with probability $p_{i}=t_{i}/T$, where $t_{i}$ is the series’ time points and $T$ is the corpus’ total time points. In addition, we follow the approach used in Moirai (Woo et al., 2024) and Timer (Liu et al., 2024) by capping the sampling probability at 0.05 to ensure a more balanced contribution from each dataset. We then randomly selected a segment from each chosen sample.
Evaluation Details. We evaluate the model on a randomly selected $10$% subset of the test data every $10^{3}$ steps to reduce computational costs. For performance measurement, we observed that non-normalized metrics like MAE and MSE are highly sensitive to the amplitude of time series data, often causing the overall average to be disproportionately influenced by high-amplitude datasets. To mitigate this issue, we primarily use the normalized metric, mean absolute percentage error (MAPE), along with the negative log-likelihood (NLL), to assess forecasting performance. For a more comprehensive understanding of TSFM scaling laws, we also include additional results using symmetric mean absolute percentage error (SMAPE), mean absolute scaled error (MASE), and continuous ranked probability score (CRPS) in Appendix D.4. Detailed descriptions of these metrics are provided in the Appendix C.3.
@@ 3 Scaling Laws for Time Series Foundation Models
In this section, we first present experimental results using the encoder-only Transformer to explore scaling laws across different data distributions. Following this, we conduct a comparative study on the scaling behavior of encoder-only and decoder-only TSFMs, Chronos and Moirai, to investigate how various scaling factors influence the characteristics of time series models.
@@ 3.1 Scaling Laws Across Data Distributions
Figure 2: Parameter Scaling. The scaling effect of total trainable model parameters on the in-distribution (ID) and out-of-distribution (OOD) forecasting performance, which is evaluated using NLL and MAPE metrics. When evaluated with NLL, both ID and OOD results follow an approximate power law scaling with parameter count, exhibiting consistent trends across different data distributions. The blue and red horizontal dashed lines represent the baselines of the exponential smoothing (ETS) forecasting method.
Parameter Scaling. In Figure 2, we display the ID and OOD performance of a wide variety of encoder-only Transformers, ranging from small models with 1K parameters through large models with 100M parameters. We trained models on the full pre-training corpus to convergence and report the minimum NLL and MAPE. We can see that both ID and OOD performance roughly follow power-law behavior over five orders of magnitude in model sizes. Formally, the power law can be expressed as:
@@
where $L$ is the performance metric function (i.e., MAPE, or NLL), $N$ is a given parameter count, $N_{c}$ is the normalization coefficient, and $\alpha_{N}$ is the exponent value that indicates the degree of performance improvement expected as we scale up $N$.
Observing the NLL metric, the lines fitting the scaling laws for both ID and OOD data exhibits a roughly constant shift and close slopes. This implies that while models incur a consistent performance bias when transferred to OOD data, their scaling patterns correlate well with their performance on the ID data. When evaluated using MAPE, the power-law for OOD scenario shows a bigger exponent value than ID scenario. This indicates that increasing model size yields greater improvements in OOD performance than ID performance. In other words, for models with weak OOD generalization capabilities, increasing model size may enable them to perform equally well on both ID and OOD data.
To evaluate whether the benefits of large-scale pre-training are warranted, we compare the pre-trained models with the classical exponential smoothing (ETS) forecasting method. The results indicate that the pre-trained models consistently outperform ETS on ID data and progressively excel on OOD data as the model size increases. This suggests that pre-trained models must reach a certain scale, at least 3M parameters in this case, to demonstrate a level of superiority on OOD data that justifies their high pre-training cost.
Figure 3: Compute Scaling. The computation scaling results indicate that model performance scales approximately according to a power law with increasing compute, consistent across both ID and OOD scenarios. The ID and OOD results illustrate that there is an lower bound for loss and MAPE on both test data under a given computational budget.
Compute Scaling. Following the similar method in (Kaplan et al., 2020), we estimate the compute budget using the formula $C=6NBS$, where $B$ is the batch size, $S$ is the number of parameter updates, i.e. the input sequence length, and 6 is the factor to account for the forward and backward passes. The ID and OOD test loss for compute budget varying over six orders of magnitude are shown in Figure 3. We see that the optimal results for each compute budget are achieved by different model sizes $N$, but the lowest loss decreases according to a approximate power law with respect to the amount of training compute. The lowest losses appear as the heavy lines, which can be fit with
@@
We observed significant noise in NLL and MAPE during training, which may be caused by the learning rate scheduler and the random sampling evaluation strategy. Similar to parameter scaling, the ID and OOD NLL show similar scaling patterns; however, when evaluated using MAPE, the OOD power law shows larger exponent values than the ID power law.
Figure 4: Data Scaling. The blue and red plots illustrate how data volume affects the ID and OOD forecasting performance of encoder-only Transformers, evaluated using NLL and MAPE metrics. The results indicate that in both scenarios, model performance scales approximately as a power law with data volume.
Data Scaling. We display empirical trends for the performance as function of dataset size $D$ in Figure 4. For the trend, we trained multiple 1B encoder-only Transformers on a series of subsets of the pre-training dataset and report the averaged evaluation results during training. We see that the NLL and MAPE can be fit well with simple power-law
@@
Different from the parameter scaling and computational scaling, we found that when evaluated using NLL, ID and OOD performance do not exhibit the same scaling behavior. Instead, ID performance is more sensitive to the scaling of dataset size compared to OOD performance. However, similar to the observations on the other two factors, the impact of data scaling on MAPE in OOD data is greater than its impact on MAPE in ID data. This suggests that the scaling of various factors yields greater improvements in OOD performance than ID performance.
Cross-distribution Scaling Effects. We summarize the key findings on how model ID and OOD performance scales with the model parameters, data volume, and compute. (1) In both ID and OOD uni-variate time series forecasting, model performance follows a simple power law as a function of model parameters, data volume, and compute. (2) For NLL, the model exhibits similar scaling patterns in both ID and OOD scenarios, in terms of model size or compute resources. (3) When using MAPE as the metric, scaling of all three factors results in greater improvements in OOD performance compared to ID performance.
Figure 5: Scaling Laws of Encoder-only vs. Decoder-only Transformer. This figure presents a comparison of scaling behaviors on NLL between encoder-only and decoder-only Transformer across three different axes: number of parameters, compute, and dataset size. Overall, both models exhibit similar scalability patterns with respect to model parameters, computation and dataset sizes across ID and OOD data, but differ in ID performance.